Suppose that S = ¥180/$. The annualized risk-free rates are 4.6% and 2.75% in Japan and the U.S., respectively. Find the expected currency movement over the next 3 years. Do not write any symbol. Express your answers as a percentage. Make sure to round your answers to the nearest 100th decimal points.
Answer #1

The Forward rate =Spot Rate*(1+Japanese Rate)^3/(1+US rate)^3 =180*(1+2.75%)^3/(1+4.6%)^3 =170.61725

The percentage change in currency rates =(170.61725-180)/180 =-5.21%